26. Consider the following information for a particular economy.
Total population = 100 million Number of employed = 15 million
Total labor force = 45 million Natural rates of unemployment = 10%

a. Find the total unemployment rate
b. Calculate the cyclical unemployment rate
1
Expert's answer
2020-10-07T08:54:58-0400

The following formula is used to calculate the unemployment rate;

"Unemployment\ rate=\dfrac{
unemployed \ persons}
{ total\ labor\ force}*100"

Total labor force = 45million

Employed people=15million

Unemployed persons="45-15=30million"


"Unemployment\ rate =\dfrac{30000000}{45000000}*100"


"=\dfrac{2}{3}*100"

"Unemployment\ rate=66.7\%"


b. Cyclical unemployment rate.

"Cyclical\ unemployment\ rate= current\ unemployment\ rate- natural\ rate\ of\ unemployment" Therefore,

"Cyclical\ unemployment\ rate= 66.7-10=55.7\%"